OpenStudy (anonymous):

Medal A soup recipe calls for 1 1/3 of a cup of milk, but the measuring cup being used is in metric. If it is known that 1 cup is equal to 8 fl. oz., find the equivalent number of mL of milk required by the recipe. Round to the nearest mL. ? mL

OpenStudy (sleepyjess):

1 cup = 8 fl. oz you need 1 1/3 cups so what is 1/3 of 8?

OpenStudy (anonymous):

yes

OpenStudy (sleepyjess):

Then once you have 1/3 of 8, add 8 ounces to that and you have 1 1/3 cups

OpenStudy (anonymous):

@sleepyjess 8/3 + 1 = 11/3 (total) Then? :D

OpenStudy (sleepyjess):

Yes

OpenStudy (anonymous):

ya what do we need to multiply

OpenStudy (sleepyjess):

Then round and you have the answer.

OpenStudy (sleepyjess):

Do as @Abhilash11 showed above. 8/3

OpenStudy (anonymous):

See @harz360 I fl = 29.5 ml So here, you actually have to multiply 29.5 to 11/3 to get the answer in ml :-)

OpenStudy (sleepyjess):

Why would you do 29.5?

OpenStudy (anonymous):

108.166

OpenStudy (anonymous):

Cause the answer required should be in ml @sleepyjess ? If that's the requirement of the question^ otherwise not required :-)

OpenStudy (sleepyjess):

OH OK I see what you are saying.

OpenStudy (anonymous):

Yeah. @harz360 Absolutley correct!

OpenStudy (anonymous):

so do we need to round that

OpenStudy (sleepyjess):

Yes, what would that round to?

OpenStudy (anonymous):

Yeah well you can. But you know what I'd suggest you take 1 fl = 30 ml though. Why cause you have 3 in the denominator, so calculation wont be a problem you know. You could cancel it directly :-)

OpenStudy (anonymous):

108.2

OpenStudy (fibonaccichick666):

if you would have learned how to set up the fractions in the last problem you could do this easily....

OpenStudy (sleepyjess):

Round to the nearest mL

OpenStudy (anonymous):

Cause 29.5 and 30 dont make that a difference. only rounding off you know :-) so 11/3 * 30 = 110 ml :-D (the answer)

OpenStudy (fibonaccichick666):

or the one before actually

OpenStudy (anonymous):

108ml

OpenStudy (anonymous):

Okay! :-)

OpenStudy (anonymous):

so the answer is 108mL

OpenStudy (anonymous):

Yupp can say so too. Depends on thr rounding off. 108 is more accurate than 110. so yeah 108

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

But then when you solve right? Take 1 fl = 30 ml. that way you can escape the multiplication part, and solve as @FibonacciChick666 said earlier, as fractions. Final equation being, 11/3 * 30

OpenStudy (anonymous):

Sure. No problem :-)

OpenStudy (anonymous):

110

OpenStudy (anonymous):

so the answer is 110 Ml or 108ML

OpenStudy (anonymous):

Aye. Take 108 :-P Accurate! and since we've solved it that way . 108! Final

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

No problem :-)